# Revealing the building blocks of tree balance: fundamental units of the Sackin and Colless Indices

### v1

@@ 4 Discussion
We conclude by noting that concerning $\Delta_{CS}$, whenever the rankings of two trees $T_{1}$ and $T_{2}$ are different between $S$ and $C$, i.e., if we have $S(T_{1})>S(T_{2})$ but $C(T_{1})<C(T_{2})$, then $\Delta_{CS}(T_{1})=C(T_{1})-S(T_{1})<C(T_{2})-S(T_{2})=\Delta_{CS}(T_{2})$. Hence, in such cases, $\Delta_{CS}$ (and thus $N_{b}$) will agree with the Colless index and not with the Sackin index concerning which tree is more balanced. Thus, it seems that for the difference $\Delta_{CS}=C-S$, indeed the Colless index is more dominant than the Sackin index (which can also be seen by Theorem 5, as it was shown there that a Sackin minimum which is not Colless minimal cannot minimize $\Delta_{CS}$ or maximize $N_{b}$, whereas all Colless minima are guaranteed to do so). An interesting question arising from our manuscript is that of finding a direct, non-recursive formula for $a(n)$, the number of $\Delta_{CS}$ minima (and $N_{b}$ and $\Delta_{SC}$ maxima). We leave this problem open for future research.
Figure 6: Ranking comparisons of the indices $N_{a}$, $S$, $C$, $\Delta_{CS}$ (whose induced rankings are the same as those of $\Delta_{SC}$ and $N_{b}$) and stairs2. Note that the depicted plots were generated with the computer algebra system Mathematica [14]. The $y$-axis contains all trees ordered according to the so-called Furnas rank [12, 11, 19], starting with the maximally balanced tree and proceeding to the caterpillar. The blue line corresponds to the caterpillar, the red line corresponds to the maximally balanced tree, which in case $n=7,8,9$ also coincides with the gfb tree. For $n=7,8$ the echelon tree also coincides with the maximally balanced tree, whereas it differs from it – and is thus depicted with its own magenta line – for $n=9$. Note that the stairs2 ranking was inverted (by using st2-st2 instead of $st2$) in order to make it an imbalance index as the others in the graphic for comparability. It can clearly be seen that $\Delta_{CS}=C-S$ is the least resolved index. However, it can also be seen that $\Delta_{CS}$ often ranks tree equally which are ranked differently between Colless on the one hand and stairs2 on the other hand – see, for instance, the light-blue line in the figure for $n=7$ as an extreme example.
@@ Acknowledgment
$=\min\limits_{j\in\{1,\ldots,n\}}\left.\begin{cases}d_{\frac{2n+2-j}{2}}+d_{\frac{2n-j}{2}}-2\left(\frac{2n-j}{2}\right)+2d_{\frac{j}{2}}-2\frac{j}{2}-2j&\text{for $j$ even}\\ 2d_{\frac{2n+1-j}{2}}-2\left(\frac{2n+1-j}{2}\right)+d_{\frac{j+1}{2}}+d_{\frac{j-1}{2}}-2\left(\frac{j-1}{2}\right)-2j&\text{for $j$ odd}\end{cases}\right\}$
$=\min\limits_{j\in\{1,\ldots,n\}}\left.\begin{cases}d_{n-\frac{j-2}{2}}+d_{n-\frac{j}{2}}-2n+2\frac{j}{2}+2d_{\frac{j}{2}}-2\frac{j}{2}-2j&\text{for $j$ even}\\ d_{n-\frac{j-1}{2}}+\underbrace{d_{n-\frac{j-1}{2}}}_{=d_{n-\frac{j+1}{2}+1}}-2n+2(\frac{j-1}{2})+d_{\frac{j+1}{2}}+d_{\frac{j-1}{2}}-2\left(\frac{j-1}{2}\right)-2j&\text{for $j$ odd}\end{cases}\right\}$
@@
    • Next, if $n$ is even, we have $\left\lfloor\frac{n}{2}\right\rfloor=\left\lfloor\frac{n+1}{2}\right\rfloor$. Then, as $1\leq\frac{j}{2}\leq\left\lfloor\frac{n+1}{2}\right\rfloor=\left\lfloor\frac{n}{2}\right\rfloor$, we can apply Equation (4) to $E$ to see that $E\geq d_{n}$.
      If, however, $n$ is odd, we have $j\leq n-1$ (as $j\leq n$, where $n$ is odd and $j$ is even), and thus $\frac{j}{2}\leq\frac{n-1}{2}\leq\left\lfloor\frac{n}{2}\right\rfloor$. Therefore, we can again apply Equation (4) to conclude once more that in this case, too, we have $E\geq d_{n}$.
  • We continue with terms $F$ and $G$, meaning we now consider the case in which $j$ is odd. We distinguish several cases.
    • If $j=1$, we have $F=d_{n}+d_{1}-2$, which, by Equation (8) is at least $d_{n+1}$ (which can be seen using $i=1$). Thus, in this case we have $F\geq d_{n+1}$. Moreover, if $j=1$, we get $G=d_{n}+d_{0}-2\cdot 0=d_{n}$, which immediately shows $G\geq d_{n}$ as desired.
    • If $j>1$ and $j$ is odd, we have $1\leq\frac{j-1}{2}<\frac{j}{2}<\frac{j+1}{2}\leq\frac{n+1}{2}$. Now, if $n$ is even, we have $j\leq n-1$ (as $j$ is odd) and $\frac{n}{2}=\left\lfloor\frac{n+1}{2}\right\rfloor$, implying $\frac{j+1}{2}=\frac{j}{2}+\frac{1}{2}\leq\frac{n-1}{2}+\frac{1}{2}=\frac{n}{2}=\lfloor\frac{n+1}{2}\rfloor$. Therefore, we can apply Equation (8) using $i=\frac{j+1}{2}$ to conclude that $F\geq d_{n+1}$. In this case, by $\frac{j-1}{2}<\frac{j+1}{2}\leq\left\lfloor\frac{n+1}{2}\right\rfloor$, we can also conclude that $\frac{j-1}{2}\leq\left\lfloor\frac{n+1}{2}\right\rfloor-1=\left\lfloor\frac{n}{2}\right\rfloor$. Thus, we can apply Equation (4) using $i=\frac{j-1}{2}$ to conclude that $G\geq d_{n}$.
      If, however, $n$ is odd, $n+1$ is even, so we have $\frac{n+1}{2}=\left\lfloor\frac{n+1}{2}\right\rfloor$ and thus, as $\frac{j+1}{2}\leq\frac{n+1}{2}=\left\lfloor\frac{n+1}{2}\right\rfloor$, we can apply Equation (8) using $i=\frac{j+1}{2}$ to $F$ and conclude $F\geq d_{n+1}$. Moreover, in this case, by $\frac{j-1}{2}<\frac{n+1}{2}$, we conclude $\frac{j-1}{2}\leq\frac{n+1}{2}-1=\frac{n-1}{2}=\left\lfloor\frac{n}{2}\right\rfloor$ (as both values of the inequality are integers as $j$ and $n$ are odd), where the last equality is true again because $n$ is odd. Since we have $\frac{j-1}{2}\leq\left\lfloor\frac{n}{2}\right\rfloor$, we can apply Equation (4) using $i=\frac{j-1}{2}$ to $G$ to conclude that $G\geq d_{n}$.

### v2

@@ 4 Discussion
We conclude by noting that concerning $\Delta_{CS}$, whenever the rankings of two trees $T_{1}$ and $T_{2}$ are different between $S$ and $C$, i.e., if we have $S(T_{1})>S(T_{2})$ but $C(T_{1})<C(T_{2})$, then $\Delta_{CS}(T_{1})=C(T_{1})-S(T_{1})<C(T_{2})-S(T_{2})=\Delta_{CS}(T_{2})$. Hence, in such cases, $\Delta_{CS}$ (and thus $N_{b}$) will agree with the Colless index and not with the Sackin index concerning which tree is more balanced. Thus, it seems that for the difference $\Delta_{CS}=C-S$, indeed the Colless index is more dominant than the Sackin index (which can also be seen by Theorem 6, as it was shown there that a Sackin minimum which is not Colless minimal cannot minimize $\Delta_{CS}$ or maximize $N_{b}$, whereas all Colless minima are guaranteed to do so). An interesting question arising from our manuscript is that of finding a direct, non-recursive formula for $a(n)$, the number of $\Delta_{CS}$ minima (and $N_{b}$ and $\Delta_{SC}$ maxima). We leave this problem open for future research. Last but not least, we also added to the literature by investigating the echelon tree more in-depth, which is known particularly from the context of the stairs2 and $B_{1}$ indices. We hope that the explicit definition presented in this manuscript will facilitate future research concerning this tree.
Figure 6: Ranking comparisons of the indices $N_{a}$, $S$, $C$, $\Delta_{CS}$ (whose induced rankings are the same as those of $\Delta_{SC}$ and $N_{b}$) and stairs2. Note that the depicted plots were generated with the computer algebra system Mathematica [14]. The $y$-axis contains all trees ordered according to the so-called Furnas rank [12, 10, 19], starting with the maximally balanced tree and proceeding to the caterpillar. The blue line corresponds to the caterpillar, the red line corresponds to the maximally balanced tree, which in case $n=7,8,9$ also coincides with the gfb tree. For $n=7,8$ the echelon tree also coincides with the maximally balanced tree, whereas it differs from it – and is thus depicted with its own magenta line – for $n=9$. Note that the stairs2 ranking was inverted (by using st2-st2 instead of $st2$) in order to make it an imbalance index as the others in the graphic for comparability. It can clearly be seen that $\Delta_{CS}=C-S$ is the least resolved index. However, it can also be seen that $\Delta_{CS}$ often ranks trees equally which are ranked differently between Colless on the one hand and stairs2 on the other hand – see, for instance, the light-blue line in the figure for $n=7$ as an extreme example.
@@ Acknowledgment
    • Next, if $n$ is even, we have $1\leq\frac{j}{2}\leq\left\lfloor\frac{n+1}{2}\right\rfloor=\left\lfloor\frac{n}{2}\right\rfloor$.
      If, however, $n$ is odd, we have $j\leq n-1$ (as $j\leq n$, where $n$ is odd and $j$ is even), and thus $\frac{j}{2}\leq\frac{n-1}{2}\leq\left\lfloor\frac{n}{2}\right\rfloor$. Therefore, in both cases we can apply Equation (4) to conclude that $E\geq d_{n}$.
  • We continue with terms $F$ and $G$, meaning we now consider the case in which $j$ is odd. We distinguish two cases.
    • If $j=1$, we have $F=d_{n}+d_{1}-2$, which, by Equation (7) is at least $d_{n+1}$ (which can be seen using $i=1$). Thus, in this case we have $F\geq d_{n+1}$. Moreover, if $j=1$, we immediately get $G=d_{n}$.
    • If $j>1$ and $j$ is odd, we have $1<\frac{j}{2}<\frac{j+1}{2}\leq\frac{n+1}{2}$. Furthermore, we either have $j\leq n-1$ if $n$ is even or $\frac{n+1}{2}=\left\lfloor\frac{n+1}{2}\right\rfloor$ if $n$ is odd. In both cases $1\leq\frac{j+1}{2}\leq\lfloor\frac{n+1}{2}\rfloor$ and we can apply Equation (7) using $i=\frac{j+1}{2}$ to conclude that $F\geq d_{n+1}$. The fact that $\frac{j+1}{2}\leq\lfloor\frac{n+1}{2}\rfloor$ also gives us $1\leq\frac{j-1}{2}=\frac{j+1}{2}-1\leq\lfloor\frac{n+1}{2}\rfloor-1\leq\lfloor\frac{n}{2}\rfloor.$ Therefore, we can apply Equation (4) using $i=\frac{j-1}{2}$ to $G$ to conclude that $G\geq d_{n}$.